# Monte Carlo Tree Diffusion for System 2 Planning

### v3

@@ A.7.7 Replanning is Necessary to Solve Complex Environments
In this subsection, we investigate the failure modes of diffusion-based planners, particularly Diffusion Forcing without replanning, in solving the visual pointmaze task. Figure 7 demonstrates how the absence of iterative corrections leads to trajectory collapse, revealing critical challenges in pixel-level planning. These findings indirectly highlight the necessity of replanning for robust trajectory generation in complex and partially observable environments.
@@ Appendix B MCTD Architecture Implementation
The Monte Carlo Tree Diffusion (MCTD) framework relies on three key elements within its tree‐search loop: sub‐plan noise scheduling, causal semi‐autoregressive denoising, and partial, jumpy denoising. First, rather than assigning a single noise schedule to the entire trajectory, each sub‐trajectory segment receives its own noise‐level schedule. This design allows different segments to transition from high to low noise at distinct rates, thereby enabling focused refinement of earlier parts of the plan while preserving flexibility in later segments.
Next, we adopt a causal noise schedule (diffusion_forcing), which renders the denoising process semi‐autoregressive: earlier plan segments undergo more thorough denoising, while subsequent ones shift more gradually from higher to lower noise. This preserves the globally consistent trajectory generation characteristic of diffusion models but also ensures that critical decisions near the beginning of the plan receive higher‐fidelity refinements. Finally, both planning and simulation employ partial and jumpy denoising via DDIM sampling (ddim), allowing the system to skip directly from noise level $t$ to $t-\Delta$. This skip‐level update accelerates the evaluation of candidate trajectories during tree expansion and supports partial denoising steps essential for exploration.
We implement these strategies with a Transformer‐based model that processes an expandable token sequence, thus avoiding any architectural constraint on the planning horizon. Each sub‐plan is annotated with a distinct noise‐level index, and the Transformer predicts the fully denoised data $x_{0}$ from the corresponding noisy tokens. Concretely, we employ an $x_{0}$‐parameterization by training the network to minimize the squared error between the predicted $\hat{x}_{0}$ and the ground‐truth $x_{0}$:
@@
Because each sub‐plan has its own noise schedule, the Transformer naturally accommodates varying horizon lengths and partial denoising intervals without requiring further modifications to the model architecture.
By integrating causal scheduling with skip‐level DDIM updates, MCTD achieves a semi‐autoregressive, partial denoising process within a tree‐search loop. Early plan segments are refined immediately, while future segments remain at higher noise until exploration or heuristic signals indicate their importance. This arrangement allows for fast approximate rollouts by skipping multiple noise levels when simulating a plan. Consequently, MCTD integrates the holistic generative capacity of diffusion models with the adaptive, branch‐and‐bound nature of Monte Carlo Tree Search, thereby enabling iterative trajectory refinement in tasks with long horizons and sparse rewards.
@@ Appendix C Detailed Comparison between Monte Carlo Guidance (MCG) and Monte Carlo Tree Diffusion (MCTD)
While MCG (Monte Carlo Guidance) utilizes Monte Carlo sampling to adjust guidance levels during denoising, it lacks the structured search capabilities of MCTD (Monte Carlo Tree Diffusion) and does not explicitly refine trajectories over time. MCG focuses on adjusting guidance strengths rather than performing an explicit trajectory search, making it less effective for long-horizon planning where iterative improvement is crucial. Additionally, MCG does not implement the four standard MCTS steps (Selection, Expansion, Simulation, and Backpropagation), limiting its ability to efficiently allocate test-time compute (TTC) for progressive refinement. Without a structured search process, MCG cannot revisit and optimize subplans, making it less scalable for complex decision-making tasks. Furthermore, MCG lacks meta-actions for adaptive exploration-exploitation, passively adjusting guidance instead of dynamically balancing search strategies. Finally, it does not incorporate fast jumpy denoising for simulation, making trajectory evaluation less efficient compared to MCTD, which efficiently estimates plan quality without costly forward model rollouts.
Appendix D Algorithms

### v5

@@ A.13.7 The Necessity of Replanning in Complex Environments
In this section, we analyze the failure modes of diffusion-based planners that do not employ replanning, such as Diffusion Forcing, in the visual pointmaze task. As illustrated in Figure 7, the absence of iterative correction leads to trajectory collapse, revealing significant challenges in pixel-based planning. These findings underscore the necessity of replanning for generating robust trajectories in complex and partially observable environments.
@@ Appendix B MCTD Architecture Implementation
The Monte Carlo Tree Diffusion (MCTD) framework relies on three key elements within its tree-search loop: sub-plan noise scheduling, causal semi-autoregressive denoising, and partial, jumpy denoising. First, rather than assigning a single noise schedule to the entire trajectory, each sub-trajectory segment receives its own noise-level schedule. This design allows different segments to transition from high to low noise at distinct rates, thereby enabling focused refinement of earlier parts of the plan while preserving flexibility in later segments.
Next, we adopt a causal noise schedule (Chen et al., 2024a), which renders the denoising process semi-autoregressive: earlier plan segments undergo more thorough denoising, while subsequent ones shift more gradually from higher to lower noise. This preserves the globally consistent trajectory generation characteristic of diffusion models but also ensures that critical decisions near the beginning of the plan receive higher-fidelity refinements. Finally, both planning and simulation employ partial and jumpy denoising via DDIM sampling (Song et al., 2021a), allowing the system to skip directly from noise level $t$ to $t-\Delta$. This skip-level update accelerates the evaluation of candidate trajectories during tree expansion and supports partial denoising steps essential for exploration.
We implement these strategies with a Transformer-based model that processes an expandable token sequence, thus avoiding any architectural constraint on the planning horizon. Each sub-plan is annotated with a distinct noise-level index, and the Transformer predicts the fully denoised data $x_{0}$ from the corresponding noisy tokens. Concretely, we employ an $x_{0}$-parameterization by training the network to minimize the squared error between the predicted $\hat{x}_{0}$ and the ground-truth $x_{0}$:
@@
Because each sub-plan has its own noise schedule, the Transformer naturally accommodates varying horizon lengths and partial denoising intervals without requiring further modifications to the model architecture.
By integrating causal scheduling with skip-level DDIM updates, MCTD achieves a semi-autoregressive, partial denoising process within a tree-search loop. Early plan segments are refined immediately, while future segments remain at higher noise until exploration or heuristic signals indicate their importance. This arrangement allows for fast approximate rollouts by skipping multiple noise levels when simulating a plan. Consequently, MCTD integrates the holistic generative capacity of diffusion models with the adaptive, branch-and-bound nature of Monte Carlo Tree Search, thereby enabling iterative trajectory refinement in tasks with long horizons and sparse rewards.
@@ Appendix C Detailed Comparison between Monte Carlo Guidance (MCG) and Monte Carlo Tree Diffusion (MCTD)
While MCG (Monte Carlo Guidance) utilizes Monte Carlo sampling to adjust guidance levels during denoising, it lacks the structured search capabilities of MCTD (Monte Carlo Tree Diffusion) and does not explicitly refine trajectories over time. MCG focuses on adjusting guidance strengths rather than performing an explicit trajectory search, making it less effective for long-horizon planning where iterative improvement is crucial. Additionally, MCG does not implement the four standard MCTS steps (Selection, Expansion, Simulation, and Backpropagation), limiting its ability to efficiently allocate inference-time scaling for progressive refinement. Without a structured search process, MCG cannot revisit and optimize subplans, making it less scalable for complex decision-making tasks. Furthermore, MCG lacks meta-actions for adaptive exploration-exploitation, passively adjusting guidance instead of dynamically balancing search strategies. Finally, it does not incorporate fast jumpy denoising for simulation, making trajectory evaluation less efficient compared to MCTD, which efficiently estimates plan quality without costly forward model rollouts.
Appendix D Additional Experiment Results
D.1 UCT Hyperparameter Ablation Study
In addition to the meta-action ablation study (Section 5.6.2), we empirically analyze the exploration-exploitation balance by varying the UCT (Kocsis & Szepesvári, 2006) hyperparameter $W$ in the formula: $v_{\text{UCT}}=v_{i}+W\sqrt{\frac{\ln N}{n_{i}}}$, where $v_{i}$ is the node’s estimated value, and $N$ and $n_{i}$ are the visit counts of the parent node and node itself, respectively.
Table 14: UCT Hyperparameter Ablation Study Results on PointMaze.
$W$ 0 (Greedy) $\sqrt{2}$ (Default) 3 5 10
Medium Success Rate (%) 88$\pm$13 100$\pm$0 100$\pm$0 98$\pm$6 98$\pm$6
Run Time (sec.) 15$\pm$1 31$\pm$4 63$\pm$10 76$\pm$23 92$\pm$55
The number of Search 105$\pm$84 77$\pm$29 94$\pm$8 120$\pm$27 134$\pm$23
Large Success Rate (%) 90$\pm$10 98$\pm$6 100$\pm$0 98$\pm$6 100$\pm$0
Run Time (sec.) 16$\pm$0 74$\pm$34 90$\pm$10 102$\pm$14 104$\pm$18
The number of Search 117$\pm$78 174$\pm$90 211$\pm$26 257$\pm$41 265$\pm$43
Giant Success Rate (%) 82$\pm$14 100$\pm$0 98$\pm$6 100$\pm$0 100$\pm$0
Run Time (sec.) 25$\pm$1 215$\pm$23 216$\pm$12 225$\pm$22 230$\pm$10
The number of Search 228$\pm$69 394$\pm$152 442$\pm$13 464$\pm$13 493$\pm$7
As shown in Table 14, with $W=0$ (pure greedy search), MCTD achieves faster inference times and requires fewer searches, but at the cost of reduced performance. This occurs because greedy search explores the tree depth-wise rapidly, reducing computational overhead of jump denoising but sacrificing thorough exploration. The default value ($W=1.141$) strikes an effective balance, while higher values increase computational costs.
D.2 Jumpiness Scale Ablation Study
Fast denoising (Song et al., 2021a) is a key concept in MCTD that enables tree search within the denoising process. The jumpiness scale $C$ determines how many denoising steps are skipped during this process. Larger $C$ values result in fewer denoising steps, trading accuracy for computational speed. We investigated how different jumpiness scales affect performance and efficiency, including a one-shot variant that denoises trajectories in a single step. The results demonstrate that proper selection of $C$ significantly impacts both model performance and efficiency.
Table 15: Jumpiness Scale Ablation Study Results on PointMaze.
$C$ 1 5 10 (Default) 20 50 One-shot
Medium Success Rate (%) 100$\pm$0 100$\pm$0 98$\pm$6 100$\pm$0 100$\pm$0 100$\pm$0
Run Time (sec.) 84$\pm$26 39$\pm$16 34$\pm$13 29$\pm$12 27$\pm$13 42$\pm$12
The number of Search 86$\pm$30 74$\pm$32 77$\pm$29 73$\pm$30 65$\pm$29 143$\pm$46
Large Success Rate (%) 100$\pm$0 100$\pm$0 100$\pm$0 98$\pm$6 94$\pm$13 63$\pm$39
Run Time (sec.) 160$\pm$69 91$\pm$35 74$\pm$34 65$\pm$32 68$\pm$49 355$\pm$183
The number of Search 176$\pm$85 188$\pm$81 174$\pm$90 167$\pm$89 174$\pm$112 301$\pm$134
Giant Success Rate (%) 100$\pm$0 98$\pm$6 100$\pm$0 100$\pm$0 88$\pm$16 10$\pm$21
Run Time (sec.) 632$\pm$239 258$\pm$115 231$\pm$103 185$\pm$88 174$\pm$76 164$\pm$2
The number of Search 390$\pm$148 379$\pm$158 394$\pm$152 363$\pm$154 387$\pm$149 500$\pm$0
For relatively simple tasks, even aggressive jumpiness (large $C$) or one-shot denoising maintains high performance. However, as task complexity increases, these approaches significantly degrade performance because value function estimation becomes less accurate under highly jumpy denoising, leading to suboptimal tree search decisions. The default setting ($C=10$) achieves an excellent balance, maintaining high success rates while reducing computational time compared to smaller jump values. Additionally, we tested one-shot decoding by directly applying predictions from a diffusion model trained to predict fully denoised data. While this represents the most computationally efficient method, it exhibits less effective search behavior in medium-sized maps and clear performance degradation in larger maps. These results demonstrate that jumpy denoising provides an optimal balance between effectiveness and efficiency.
Appendix E Algorithms